# Strain stiffening due to stretching of entangled particles in random packings of granular materials

### v1

@@ Appendix G Relation between force and plastic deformation of staples
To measure the slope $\partial F_{p}/\partial\theta_{p}$ plastic deformation for individual pairs of entangled staples, we hooked two staples together and pulled on the opposite ends of the staples to mimic the arrangement of entangled particles pulling against each other in a random packing. The pair of particles were entangled, while clamping the unhooked arms of each staple, as shown in the inset of Fig. 24. To arrange the particles so that the entangling would provide some resistance to pulling, the clamped arms were rotated 90 degrees in the horizontal plane relative to each other, and the backbones were tilted relative to vertical so the hook ends would not slide apart when initially pulled on. We then pulled on the staples moving the top clamp at a rate of 0.05 mm/s while recording the maximum extensional force $F_{p}$ applied by the materials tester. After the peak force, plastic deformation occurred, typically causing the hooked staples to slip and the force to drop suddenly. Bend angles $\theta_{p}$ were measured after the force was completely removed to record the angle of permanent plastic deformation. Each measurement point was done with a new staple to avoid effects of a history of plastic deformation. We also varied the initial backbone orientations relative to the vertical by 10 to 23 in different cases in a range where we could obtain stable entangling. In this range, the level arm distance of the point of contact typically stabilized at $l=0.6\pm 0.2$ perpendicular to the backbone due to the curvature near the corner and width of the staples. While there is a lot variation in the data from run to run, we found no clear systematic dependence of the results on this initial orientation, so we expect these tests are fairly representative of entangling in a random packing. After pulling, we removed the staples from the setup, relieving the stress, and measured the bend angle of the staples in terms of a change in bend angle between the arms and the backbone relative to the initial 90 angle. The reported angles are the mean plus or minus the standard deviation of the change of the 4 angles – 2 angles for each staple.

### v2

@@ Appendix G Relation between force and plastic deformation of staples
To measure the slope $\partial F_{p}/\partial\theta_{p}$ for plastic deformation of individual pairs of entangled staples, we entangled two staples together and pulled on the opposite ends of the staples to mimic the arrangement of entangled particles pulling against each other in a random packing. The pair of particles were hooked around each other, while clamping the unhooked arms of each staple, as shown in the inset of Fig. 24. To arrange the particles so that the entangling would provide some resistance to pulling, the clamped arms were rotated 90 degrees in the horizontal plane relative to each other, and the backbones were tilted relative to vertical so the hook ends would not slide apart when initially pulled on. We then pulled on the staples moving the top clamp at a rate of 0.05 mm/s while recording the maximum extensional force $F_{p}$ applied by the materials tester. After the peak force, plastic deformation occurred, typically causing the hooked staples to slip and the force to drop suddenly. Bend angles $\theta_{p}$ were measured in terms of a change in angle between the arms and the backbone relative to the initial 90 angle, and this was measured after the force was completely removed to record the angle of permanent plastic deformation. The reported bend angles $\theta_{p}$ are the mean of the 4 angles – 2 angles for each staple. Each measurement point was done with a new staple to avoid effects of a history of plastic deformation. We also varied the initial backbone orientations relative to the vertical by 10 to 23 in different cases in a range where we could obtain stable entangling to provide representative cases for random packings. In this range, the lever arm distance of the point of contact typically stabilized at $l=0.6\pm 0.2$ perpendicular to the backbone due to the curvature near the corner and width of the staples. While there is a lot of variation in the data from run to run, we found no clear systematic dependence of the results on this initial orientation.